OpenStudy (denise):

f(x)=-5x^2+6x-5 How do I find the realative maxima and minima

OpenStudy (anonymous):

Find derivative

OpenStudy (denise):

10x-6=0

OpenStudy (anonymous):

x=3/5

OpenStudy (denise):

x=3/5 is my minima ?

OpenStudy (anonymous):

f''(x)=-10 which is negative, so function has a relative maxima at x=3/5

OpenStudy (amistre64):

its a parabola....a big bowl turned upside down :)

OpenStudy (amistre64):

the maxima as at what is refered to as its vertex; the point where it turns around and heads back down again; the hgihest point

OpenStudy (anonymous):

Refer to the attached plot.